The Linux Sound Playing HOWTO

Yoo C. Chung, wacko@power1.snu.ac.kr v1.2, 12 November 1996 This document lists applications for Linux that play various sound formats.

Contents

1 Introduction 2 1.1 Copyright of this document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Copyright of the listed applications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Where to get this document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 Feedback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.5 Acknowledgments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Playing Various Sound Formats 3

2.1 MIDI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.1.1 The adagio package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.1.2 timidity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.1.3 playmidi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.2 Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2.1 tracker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2.2 s3mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2.3 gmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.4 mod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.5 MikMod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.6 nspmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2.7 yampmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.3 MPEG audio streams . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.3.1 maplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.3.2 Layer 3 Shareware Encoder/Decoder . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.4 WAV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4.1 wavplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4.2 XWave . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.5 Other stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.5.1 sox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.5.2 bplay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2.5.3 RealAudio Player . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2.5.4 cat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3 Other useful sound utilities 10

3.1 volume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.2 setmixer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.3 Tickle Music . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 References 11

1 Introduction

This is the Sound Playing HOWTO. It lists the many sound formats and the applications that can be used to play them. It also lists some hacks and advice on using these applications. There are also some other interesting applications related to sound not directly related to playback. However, this document does not describe how one can setup a Linux system for sound support. Refer to the Linux Sound HOWTO by Jeff Tranter for instructions on setting up a Linux system for sound support and the supported sound hardware.

This deals with normal user sound applications. That is, it is only concerned about what the average user needs to know on the application side of sound, not exotic stuff like speech synthesis, or hardware stuff which is dealt in the Sound HOWTO.

1.1 Copyright of this document

This document can be freely distributed and modified (I would appreciate it if I were notified of any modifications), as long as this copyright notice is preserved. However, it cannot be placed under any further restrictions, and a modified document must have the same copyright as this one. Also, credit must be given where due.

1.2 Copyright of the listed applications

If there is no mention of any copyright, then the application is under the GNU General Public License.

1.3 Where to get this document

This document can be obtained from the Linux Documentation Project

<http://sunsite.unc.edu/mdw/linux.html>.

A Korean version of this document (rather outdated) is available at

<http://plaza.snu.ac.kr/~wacko/howto/Sound-Playing-HOWTO.ks>.

A Japanese version of this document is available at

<http://jf.gee.kyoto-u.ac.jp/JF/JF-ftp/euc/Sound-Playing-HOWTO.euc>.

1.4 Feedback

I am not omniscient, and I don’t use all the applications in here (a few I can’t even try), so there are bound to be mistakes. Also, programs usually continuously evolve, so documentation tends to get out of date. Therefore, if you find anything wrong, please send me any corrections. Suggestions or additions to this document are welcome, too.

1.5 Acknowledgments

All the authors of the applications in this HOWTO. Also, Hannu Savolainen for the great sound driver and Linus Torvalds for the great underlying OS.

I’d also like to thank Raymond Nijssen (raymond@es.ele.tue.nl), Jeroen Rutten

(jeroen@es.ele.tue.nl), Antonio Perez (aperez@arrakis.es), and Ian Jackson

(ijackson@gnu.ai.mit.edu) for their information and help.

2 Playing Various Sound Formats

There are many kinds of sound formats (WAV, MIDI, MPEG etc.). Below, we list the various formats and the applications that can be used to play them.

2.1 MIDI

MIDI stands for Musical Instrument Device Interface. MIDI files usually have the extension .mid. They contain sequencing information, that is, information on when to play what instrument in what way, etc. Depending on your hardware (and maybe the software you use to play them), the sound might be awesome, or it might be downright crappy.

2.1.1 The adagio package

This package includes mp (a command-line MIDI file player) and xmp (an XView based MIDI file player). You will need the SlingShot extensions to use xmp. It also contains other programs for playing Adagio scores.

If you have a GUS, mp can also play MOD files (see section 2.2 for more information on modules).

One little annoying bug (as of version 0.5) is that the sound breaks at the end. Namely, instead of ending the sound the way the MIDI file specifies, it ends by playing the note right before the last one in a long interval. It hasn’t stopped me from using mp, but it might prevent someone from using it for ‘real’ work. It also starts up relatively slowly.

The package does not mention any copyright (at least none that I can find), so I assume it can be freely redistributed and modified. Don’t hold me to my word, though.

It is a port of the CMU MIDI Toolkit to Linux (though there was enough added to make this questionable)

by Greg Lee (lee@uhunix.uhcc.hawaii.edu).
It can be found at <ftp://tsx-11.mit.edu/pub/linux/packages/sound/adagio05.tar.gz>. The

binaries included here are in a.out format (linked with ancient libraries), and the xmp binary segfaults in a X11R6 environment (XFree86 3.1.1, libc 4.7.2). The mp binary works fine.

You will need a bit of hackery to compile it. Actually, it’s not much of a hackery. All you have to do is to include the -lfl switch at the end of SHROBJ and XMPOBJ in the Makefile. This is to link in the flex library, which is not linked in by default. Then follow the installation instructions. And don’t forget to have XView and the SlingShot extensions installed if you want to compile xmp.

2.1.2 timidity

Some people recommend this experimental program because of good sound quality (which is very true, it’s much better than mp on a Sound Blaster 16, though probably won’t be much different on a GUS). However, it suffers from high CPU loads. It plays MIDI by first converting MIDI to WAV and then plays the WAV (you can also convert a MIDI file to a WAV file without playing if you want). This is the reason for its CPU intensive nature.

It also has an optional ncurses, SLang, Tcl/Tk or Motif interface.

You need Gravis Ultrasound patch files to use this. Look into the FAQ included with timidity for more information.

The latest version of timidity can be found at <http://www.clinet.fi/~toivonen/timidity/>. This

page also contains a link to a small library of GUS patches. The Motif version can be found at

<http://www.loria.fr/~pagel>.

It is written by Tuukka Toivonen (titoivon@snakemail.hut.fi).

2.1.3 playmidi

This is a MIDI player that plays to FM, GUS, and external MIDI. It is supposed to have a faster startup time compared to other MIDI players. It is also able to play Creative Music Files, Microsoft RIFF files, and large MIDI archives from games such as Ultima 7.

It has an X interface and a SVGA interface. It also has an option for real time playback with tracking all the notes on each channel and the current playback clock (included automatically with xplaymidi and

splaymidi).

You should do something like

     $ splaymidi foo.mid; stty sane

if you are going to use the SVGA interface, since it doesn’t reset the terminal tty mode properly. The SVGA interface may be removed in the near future.

It is written by Nathan Laredo (laredo@gnu.ai.mit.edu or laredo@ix.netcom.com).
It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/playmidi-2.3.tar.gz>.

2.2 Modules

Modules (in computer music) are digital music files, made up of a set of samples and sequencing information, telling the player when to play which sample (instrument) on which track at what pitch, optionally performing an effect like vibrato, for example.

An advantage it has over MIDI is that it can include almost any kind of sound (including human voices). Another is that it sounds just about the same on any platform, because the samples are in the module. A disadvantage it has is that it has a much larger file size compared to MIDI. Another one is that it has no real standard format (the only ‘real’ one is the ProTracker, which many modules aren’t quite compatible with). It originated on the Amiga.

They usually have the extension .mod. There are many other extensions depending on what format they are in.

2.2.1 tracker

This very portable program (it has been ported to many platforms) plays Soundtracker and Protracker music modules. It uses 16 bit stereo output, and I consider the quality to be very good. If you need a simple way to reduce CPU load use the -mono option.

This is a giftware program (quoting the author). It is by Marc Espie (Marc.Espie@ens.fr).

A version of this with the Makefile already tweaked for Linux is at

<ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/tracker-4.3-linux.tar.gz>.

2.2.2 s3mod

This plays 4/6/8 track MOD modules and Scream Tracker 3 modules. It uses 8 bit mono output with a sampling rate of 22000 Hz by default. You can use the option -s to enable stereo, -b to enable 16 bit output, and -f to set the sampling frequency. However, the sound output is worse than tracker (some noise), so I recommend using tracker instead of s3mod for playing ordinary MOD files (unless you have an underpowered machine). It has a much smaller CPU load compared to tracker.

It is copyrighted by Daniel Marks and David Jeske (jeske@uiuc.edu), but you can do anything you want with it (except that you can’t claim you wrote it).

It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/s3mod-v1.09.tar.gz>.

2.2.3 gmod

This is a music module player for the Gravis Ultrasound card. 4/6/8 channel MOD, 8 channel 669, MultiTracker (MTM), UltraTracker (ULT), FastTracker (XM), and ScreamTracker III (S3M) are the supported formats.

It requires version 3.0 or later sound driver. And a GUS, of course. You may need to modify the kernel to make volume control work the way you want.

This has an X interface. It uses the QT toolkit (needs version 0.99 or greater). Check the QT toolkit homepage <http://www.troll.no/> for information on QT.

This can be freely distributed. It was originally written by Hannu Savolainen, and now maintained by

Andrew J. Robinson (robinson@cnj.digex.net).
It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/gmod+x-3.0.1.tgz>.

2.2.4 mod

This beta program plays MODs (15/31-instrument, up to 32 voices), MTMs, ULTs and S3Ms on the Gravis Ultrasound card. It can also use packed modules if you have gzip, lharc, unzip, and unarj installed. It cannot play Powerpacked modules or modules packed with some Amiga composers (”PACK” signature).

This requires at least version 3.0 of the sound driver. It won’t work with the 2.90-2 or earlier version of the sound driver. The text interface requires ncurses. There is also an X interface included, which uses Tcl/Tk.

It is written by Mikael Nordqvist (mech@df.lth.se or d91mn@efd.lth.se).

It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/mod-v0.81.tgz>.

2.2.5 MikMod

This portable module player plays XM, ULT, STM, S3M, MTM, MOD and UNI formats. (The UNI format is an internal format used by MikMod.) It has support for zipped module files. It uses 16 bit stereo for the sound output. Use the -m option (for mono output) if you need a simple way to lower the CPU load.

The Unix version uses ncurses for its interface.

It is written by Jean-Paul Mikkers (mikmak@via.nl). The Unix version has a lot of modifications by Steve McIntyre (sam1007@cam.ac.uk). This is shareware that has to be registered if you want to use it commercially. You also need permission to redistribute it commercially (non-commercial redistribution does not need such permission).

The Unix version can be found at

<ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/mikmod-2.13-unix.tar.gz>.

2.2.6 nspmod

This is an alpha module player which can play MTM, S3M, and MOD modules. It is intended to be a module player for soundcards without a DSP (not to be confused with what Creative Labs calls a DSP). It has a CPU load somewhat similar compared to tracker.

It has a feature which lets modules loop if they want to. The number of loops can be limited by the -l option. It uses only 8 bit sound output (as of version 0.1).

This was written by Toru Egashira (toru@jms.jeton.or.jp).

It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/nspmod-0.1.tar.gz>.

2.2.7 yampmod

This alpha program was designed to play 4-channel modules using the minimum of CPU resources. It was not designed to produce high quality sound. So the only sound output it produces is 22 kHz mono output. Also, the output isn’t as clean as it should be, reflecting its alpha status.

It was written by David Groves (djg@djghome.demon.co.uk).

It is available at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/yampmod-0.1.tar.gz>.

2.3 MPEG audio streams

MPEG is a standard specifying the coding of video and the associated audio for digital storage. MPEG is usually associated with video, but the audio part of the standard can be used separately. The audio part of the MPEG standard defines three layers, layer I, II, and III. Players that can decode higher layers can also decode lower layers (e.g. layer III players can play layer II files). Layer I MPEG audio files usually have the extension .mpg (so if there is a file with this extension that can’t be played by a MPEG video player, it’s probably an audio stream), layer II usually have .mp2, and layer III usually have .mp3. The audio compression is pretty good. A two megabyte layer II MPEG audio file will probably take up 25 megabytes for a raw PCM sample file with the same quality.

2.3.1 maplay

This MPEG audio stream player only has support for layer I and layer II streams. It lacks support for layer III as of version 1.2. It supports 16 bit sound cards on Linux (though some 16 bit sound cards are rumored not to work with maplay).

It is very CPU intensive, taking up to about 55% CPU time on a 60MHz Pentium. The output is intolerable on a 66MHz 486 because the CPU just can’t catch up with the sound. If this happens to you, try playing only one side of the audio stream (with the -l or -r option), instead of the default stereo.

A slight change in one of the files may be necessary in order to compile it. Namely, you may need to add the following line to the beginning of the file configuration.sh.

     #! /bin/sh
The author is Tobias Bading (mailto:bading@cs.tu-berlin.de).

maplay can be found at

<ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/maplay1 2.tar.gz>.

2.3.2 Layer 3 Shareware Encoder/Decoder

This is actually a converter that converts MPEG Layer 3 audio streams to WAV, AIFF, SND, AIFC, or just raw PCM sample files. The Linux version does not directly output the sound to the soundcard. One has to first convert it to some other format.

However, when you try to play a converted file using sox, you’ll probably just get noise because the word order in the PCM samples is not right (at least on Intel platforms). You need to give sox the option -x to solve this problem. But there are some players that don’t have to be told that the word order is wrong, so you might not have to worry about this.

If you have a really fast computer (probably at least a 90Mhz Pentium), then you can try to play it directly without having to first convert the MPEG audio file to another format like in the following example (this example assumes that you’re using sox and playing a 44.1 kHz stereo sample).

     $ l3dec foo.mp3 -sto | play -t raw -x -u -w -c 2 -r 44100 -

The number after -r is the sample rate of the audio stream, and the number after -c depends on whether it is mono or stereo (or even quad). If this looks too complicated, you can use something like a shell script or an alias. Or you can use mp3play (found at

<ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/mp3play.tar.gz>), which sets all the

options and plays it for you (mp3play includes l3dec and uses wavplay, which is also included).

This is shareware copyrighted by Fraunhofer-IIS. A demo version for Linux on x86 systems is available at

<ftp://ftp.fhg.de/pub/layer3/l3v261.linux.tar.gz>. The demo version only converts layer III audio

streams.

2.4 WAV

Quote from the sox man page:

These appear to be very similar to IFF files, but not the same. They are the native sound file format of Windows 3.1. Obviously, Windows 3.1 is of such incredible importance to the computer industry that it just had to have its own sound file format.

These usually have the extension .wav.

Also see section 2.5.1 and 2.5.2 for other WAV players besides the ones listed here.

2.4.1 wavplay

This program has a command line interface and an X Window interface for playing and recording in WAV format. It uses locking so that only one sound may be played at a time. Its locking capabilities can also be used separately from its sound playing capabilities.

The copyright is different for each part of the program (being created by different authors), but basically you can do whatever you want with it as long as the proper credits are given. It was originally written by Andre Fuechsel (af1@irz.inf.tu-dresden.de), and the current implementation is by Warren W. Gay

(bx249@freenet.toronto.on.ca or wwg@ica.net).
It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/wavplay021w1.tar.gz>.

2.4.2 XWave

This program is based on wavplay. It too has an X Window interface for playing and recording in WAV format. However, it also has some editing capabilities and can add effects.

This is still in a rather buggy stage, so the author would appreciate any bug reports, bugfixes, or new code.

This is registered freeware, in other words you can freely redistribute it as long as you register it by sending the author a color picture postcard. But it’s OK to use it unregistered, as long as you don’t get caught???

The author is Will Fish (fishwj@ee.port.ac.uk).

XWave can be found at <ftp://ftp.ee.port.ac.uk/pub/linux/sound/xwave0.01-1.tar.gz>.

2.5 Other stuff

This section lists stuff that play sound formats that don’t deserve a separate section (i.e. formats that have only one player available), or players that play more than one format.

2.5.1 sox

This program is actually a converter, that is, it converts one sound format to another. When invoked as play, however, it plays the sound (the play application in the Sound HOWTO probably refers to this). It supports raw (no header) binary and textual data, IRCAM Sound Files, Sound Blaster .voc, SPARC .au (w/header), Mac HCOM, PC/DOS .sou, Sndtool, and Sounder, NeXT .snd, Windows 3.1 RIFF/WAV, Turtle Beach .smp, CD-R, and Apple/SGI AIFF and 8SVX formats

Since somewhere in the 1.3.6x kernels, you might have to make a small change in one file to make it play the sound directly. Namely, you may have to change line 179 in sbdsp.c from

             if (abuf_size < 4096 || abuf_size > 65536) {

to

             if (abuf_size < 1 || abuf_size > 65536) {

But then again, you might not have to do this. But doing this won’t break anything.

It is written and copyrighted by many people, and can be used for any purpose.

It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/convert/Lsox-linux.tar.gz>.

A more recent version by Chris Bagwell (cbagwell@mwsun026.aud.alcatel.com) (based on the latest gamma version of the original sox) can be found at

<ftp://sunsite.unc.edu/pub/Linux/apps/sound/convert/sox-11gamma-cb2.tar.gz>.

2.5.2 bplay

This beta program plays raw audio, WAV, and VOC files. It’s also able to record to these files. It uses a variety of techniques to get the highest speed possible so that it can run acceptably even on slow machines. One of these techniques require that the installed programs be setuid root. The paranoid hoping to use this may want to use the Debian package by Ian Jackson (ijackson@gnu.ai.mit.edu), which disables the feature that needs the setuid bit.

The author is David Monro (davidm@gh.cs.usyd.edu.au).

It can be obtained from

<ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/bplay-0.95.tar.gz>.

2.5.3 RealAudio Player

This lets you listen to sound in real time over the Internet without downloading the whole sound file first. It could be used stand alone, but it is really intended to be used along with a web browser (the explicitly supported ones are Mosaic and Netscape). It cannot be used without X (you wouldn’t be able to get it working with Lynx in a text console).

This is by Progressive Networks, Inc. This cannot be redistributed, modified etc. Look at the license for exact details on what you can do. It can be obtained by registering with no cost at the RealAudio home

page <http://www.realaudio.com/>.

2.5.4 cat

One might think what cat, the sometimes overused concatenating utility, has to do with playing sounds. I’ll show a use of it through an example.

     $ cat sample.voc > /dev/dsp
     $ cat sample.wav > /dev/dsp
     $ cat sample.au > /dev/audio

Doing a cat of an .au file to /dev/audio will usually work, and if you’re lucky enough that the file has the correct byte order (for your platform) etc., a cat of a sound file that uses PCM samples (like .wav or .voc) to /dev/dsp might even sound right.

This isn’t a totally useless use of cat. It might be useful, for example, if you have a sound file that none of your programs recognize, and you know that it uses PCM samples, then you might be able to get a very approximate idea on how it sounds like this way (if you’re lucky).

3 Other useful sound utilities

This section has nothing to do with the actual playing of sound files. Rather, it is a collection of some sound utilities that one might find useful.

3.1 volume

This is a simple command line interface for controlling the volume (what else could it be?). It also has a separate program with a Tcl/Tk interface included in the package for controlling the volume and playing .au sound files. A very simple Tcl/Tk CD player is also included.

This is Freeware and it is written by Sam Lantinga (slouken@cs.ucdavis.edu).

It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/volume-2.1.tar.gz>.

3.2 setmixer

This is a command line mixer control program.

The author is Michal Jaegermann (michal@ellpspace.math.ualberta.ca).

3.3 Tickle Music

This beta Tcl/Tk program is a music file browser that allows you to play various sound formats as long as an appropriate program to play it is on your system. By default gmod is used for playing MOD files and mp for playing MIDI files (you can change the source to use other programs).

It is written and copyrighted by Shannon Hendrix (shendrix@pcs.cnu.edu or

shendrix@escape.widomaker.com).
It can be found at <ftp://sunsite.unc.edu/pub/Linux/apps/sound/players/tmusic-1.0.tar.gz>.

4 References

1. The documentation included with the applications in this document.

2. The Linux Sound HOWTO. It is available from the Linux Documentation Project

     <http://sunsite.unc.edu/mdw/linux.html>.
  3. The Linux MIDI and Sound Pages <http://www.digiserve.com/ar/linux-snd/>
  4. alt.binaries.sounds.mods FAQ
  5. MPEG Audio Layer 3 FAQ <http://www.iis.fhg.de/departs/amm/layer3/sw/index.html>